How many kilograms of brass containing 63% copper must be melted with 1100 of brass containing 72% copper to produce a new brass containing 67% copper?
Answer by stanbon(75887)

How many kilograms of brass containing 63% copper must be melted with 1100 of brass containing 72% copper to produce a new brass containing 67% copper?
----
Equation:
cop + cop = cop
0.63x + 0.72*1100 = 0.67(x+1100)
--------
63x + 72*1100 = 67x + 67*1100
-----
4x = 5*1100
------
x = 5*275
x = 1375 units (amt. of 63% brass needed)
--------------
Cheers,
Stan H.
---------------